2 Question 3 Fill in the blanks to write the following equation as an algebraic equation and then solve for X. Two subtracted from seven times a number is 12. I dont know One attempt You answered 2 out of 2 correctly Asking up to
Answer #1

Let's consider a number is "x"

hence from the given information:

7x-2=12

7x=12+2

7x=14

x=14/7

x=2

Therefore x=2